OpenStudy (idealist10):

Solve xy^(3)y'=y^4+x^4 explicitly.

OpenStudy (anonymous):

Another substitution like \(y=ux\): \[\begin{align*} xy^3y'&=y^4+x^4\\ y'&=\frac{y}{x}+\frac{x^3}{y^3} \end{align*}\] Letting \(u=\dfrac{y}{x}\) and using \(y'=u'x+u\), you have \[\begin{align*}u'x+u&=u+\frac{1}{u^3}\\\\ u'x&=\frac{1}{u^3}\\\\ u^3~du&=\dfrac{dx}{x}\end{align*}\]

OpenStudy (idealist10):

How did you came up with the substitution y=ux?

OpenStudy (anonymous):

\(y=ux\) is a pretty standard sub (refer to the link I posted a while ago), and it comes from being able to write the equation in terms of the expression \(\dfrac{y}{x}\). I saw that you could do that, so I used this sub.

OpenStudy (idealist10):

Okay

OpenStudy (idealist10):

After I integrate, I got u^4/4=ln abs(x)+C, and u^4=4ln abs(x)+C, u=(4*ln abs(x)+C)^(1/4) y/x=(4*ln abs(x)+C)^(1/4) y=x(4*ln abs(x)+C)^(1/4)

OpenStudy (idealist10):

Thanks!